OpenStudy (anonymous):

f(x)=3x+2 g(x)=-2x^2-3x+5 Find g(f(x)) @Jhannybean

OpenStudy (anonymous):

so first I would combine them?

OpenStudy (anonymous):

so 3x+2-2x^2-3x+5=g(f(x))

OpenStudy (anonymous):

g(f(x))=-2x^2+7 right?@Jhannybean

jimthompson5910 (jim_thompson5910):

g(f(x)) means you start with g(x), then you replace x with f(x) like this g(x)=-2x^2-3x+5 g(f(x))=-2(f(x))^2-3(f(x))+5 ... replace ALL copies of x with f(x) g(f(x))=-2(3x+2)^2-3(3x+2)+5 ... on the right side, replace ALL copies of f(x) with 3x+2 now expand and simplify

OpenStudy (anonymous):

ok so @jim_thompson5910 was @Jhannybean wrong with how she show'd me the other problems?

jimthompson5910 (jim_thompson5910):

no I didn't say that

OpenStudy (anonymous):

i know but this isn't how she showd me

jimthompson5910 (jim_thompson5910):

its possible before you were finding f(x) + g(x) or f(x) - g(x) but now you've moved onto composite functions

OpenStudy (anonymous):

so i wanna make sure i did the others correct ly

jimthompson5910 (jim_thompson5910):

hopefully you know the difference between them?

OpenStudy (anonymous):

they were all composite functions?

jimthompson5910 (jim_thompson5910):

oh they were? are you sure?

OpenStudy (anonymous):

yup, thats the entire worksheet

jimthompson5910 (jim_thompson5910):

alright, then yeah, it was gone about the wrong way

jimthompson5910 (jim_thompson5910):

use the method I showed above

OpenStudy (anonymous):

uhhh great so would f(x)=-3+2 g(x)=3x^2-x+2 Find (f+g)(x) (f+g)(x)=3x^2-4x is that right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

oh those aren't composite functions

jimthompson5910 (jim_thompson5910):

(f+g)(x) = f(x) + g(x)

jimthompson5910 (jim_thompson5910):

that's not a composite function

OpenStudy (anonymous):

what's the difference?

jimthompson5910 (jim_thompson5910):

f(g(x)) or g(f(x)) those are composite functions

OpenStudy (anonymous):

ooooooooooo

OpenStudy (anonymous):

but was it still right?

jimthompson5910 (jim_thompson5910):

the difference is that with composite functions, you were plugging one entire function into the input of another

jimthompson5910 (jim_thompson5910):

f(x)=-3+2 is that right??

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

there's no x?

OpenStudy (anonymous):

there is it's -3x+2

OpenStudy (anonymous):

my bad

jimthompson5910 (jim_thompson5910):

so f(x) = -3x+2 g(x) = 3x^2-x+2 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

find (f+g)(x)

OpenStudy (anonymous):

I got 3x^2-4x

jimthompson5910 (jim_thompson5910):

(f+g)(x) = f(x) + g(x) (f+g)(x) = -3x+2 + 3x^2-x+2 (f+g)(x) = 3x^2-4x+4

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

I got 3x^2-4x cuz don't the -2 n +2 cancel eachother out?

jimthompson5910 (jim_thompson5910):

the two '2's are positive though

jimthompson5910 (jim_thompson5910):

f(x) = -3x+2 g(x) = 3x^2-x+2

OpenStudy (anonymous):

ooooo ur right

jimthompson5910 (jim_thompson5910):

silly mistake

jimthompson5910 (jim_thompson5910):

but you pretty much got it

OpenStudy (anonymous):

haha :)

OpenStudy (anonymous):

and this one too? f(x)=x-1 g(x)=2x^2-2x-5 Find g(f(x)) g(f(x))=2x^2-6x-1 is that right?

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

u rock!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

thx lol

OpenStudy (anonymous):

well that took long

OpenStudy (anonymous):

wait so g(f(x))=-2(3x+2)^2-3(3x+2)+5 would simplify to (-6x+-4)^2-(9x+6)5 @jim_thompson5910

OpenStudy (anonymous):

@genius12 it was three problems

OpenStudy (anonymous):

ok lol

jimthompson5910 (jim_thompson5910):

not sure how you got (-6x+-4)^2-(9x+6)5

OpenStudy (anonymous):

multiply

jimthompson5910 (jim_thompson5910):

btw you do NOT distribute the -2 through to (3x+2) you have to FOIL out (3x+2)^2 first, then you can distribute

OpenStudy (anonymous):

o one sec

OpenStudy (anonymous):

(21x+4)

jimthompson5910 (jim_thompson5910):

not sure how you got (3x+2)^2 to turn into 21x+4

OpenStudy (anonymous):

-18x^2-33x-9

OpenStudy (anonymous):

it's magic, it's magic.

OpenStudy (anonymous):

I'm a lil confused

jimthompson5910 (jim_thompson5910):

what do you get when you FOIL out or expand (3x+2)^2 keep in mind that (3x+2)^2 = (3x+2)(3x+2) (3x+2)^2 = 3x(3x+2) + 2(3x+2) (3x+2)^2 = ???

OpenStudy (anonymous):

@jim_thompson5910 my final answer to the problem would be -3(2x+3)(3x+1)

OpenStudy (anonymous):

is that correct? lol btw I love how there are like 5 ppl just watching my problem

jimthompson5910 (jim_thompson5910):

it's that fascinating lol, but that's if they want you to factor but I'm assuming they just want you to write it in standard form

OpenStudy (anonymous):

yeah, I think ur right, in that case it'd be -18x^2-33x-9 i think

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

coolio ty!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

g(f(x))=-2(3x+2)^2-3(3x+2)+5 g(f(x)) = -18x^2-33x-9

jimthompson5910 (jim_thompson5910):

yw